Circular No. 8725

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 M2 (SPACEWATCH) An apparently asteroidal object discovered in the course of the Space-

watch project (discovery observation tabulated below) has been found to be cometary following posting on the ‘NEOCP’. J. Young reports that stacked Table Mountain exposures from June 21.23–21.28 UT show a circular 6′′ coma with little if any condensation and no tail, while his images from June 20 showed only a stellar appearance; on June 22.2, in better seeing, Young’s frames show what appears to be a varying degree of condensation, suggesting some rotation, with a 4′′–6′′ coma of mag 20.6. R. S. McMillan reports that CCD frames taken with the 1.8-m f/2.7 Spacewatch II reflector on June 22.35 show the object to be possibly more diffuse than surrounding stars, but seeing was described as ‘soft’.

2006 UT  α_{2000}  δ_{2000}  Mag.
June 19.31525  16^{h}03^{m}44.^{s}52  +2^{o}13′53′′.0  21.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-M27.

T = 2005 Nov. 20.441 TT  ω = 115.^{o}255
Ω = 27.083 }2000.0
q = 5.20585 AU  i = 150.300

SUPERNOVAE 2006de, 2006df, 2006dg, 2006dh Four additional supernovae have been discovered on unfiltered CCD

images: 2006df by the ‘Nearby Supernova Factory’ collaboration (cf. CBET 556), and the rest by N. Lee and W. Li (LOSS/KAIT; cf. IAUC 8723).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006de June 2.49 21^{h}35^{m}33.^{s}33 −19^{o}01′46′′.5 17.2 11′′.2 W, 4′′.6 N
2006df June 16.5 21 49 57.67 +32 01 58.7 19.5  –
2006dg June 17.47 23 45 54.38 +12 03 59.0 18.3 10′′.7 W, 17′′.1 N
2006dh June 21.25 13 41 43.08 +40 52 36.8 18.4 9′′.6 E, 8′′.2 N

Additional KAIT magnitudes for 2006de in MCG –03-55-3: 2005 Oct. 22.19 UT, [19.0; 2006 June 11.47, 17.5. SN 2006df is a type-IIn supernova. Additional KAIT magnitudes for 2006dg in IC 1508: 2005 Nov. 21.19, [19.5; 2006 June 22.46, 18.5. Additional KAIT magnitudes for 2006dh in UGC 8670: May 18.34, [19.2; June 4.30, [18.2; 22.22, 18.3.

2006 June 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green